OpenStudy (anonymous):

3x-8(3-x)=53 show work

OpenStudy (anonymous):

what do you want to know about this equation?

OpenStudy (anonymous):

(3x-8)(3-x) just multiply it out as you would a normal multipication: a) -x * -8 = +8x b)-x * 3x = -3x^2 c) 3 * -8 = -16 d) 3 * 3x = 9x putting this together (highest powers first) you get: -3x^2 +9x +8x -16 = 53 lets change the negative sign: 3x^2-9x-8x+16 = -53 3x^2-17x+16+53=0 3x^2-17+59=0 and use quadratic formula to solve this, to get two solutions.